OpenStudy (anonymous):

integral of square root (1-sin2x)

OpenStudy (sumi29):

Well where exactly are you stuck?

OpenStudy (anonymous):

is it \[\int \sqrt{1-\sin^2(x)}dx\]?

OpenStudy (anonymous):

or \[\int \sqrt{1-\sin(2x)}dx\]

OpenStudy (anonymous):

1-sin(2x) = sin(x)^2 + cos(x)^2 - 2sinxcosx does this help ?